# Deciphering the dynamics of nuclear collisions with elongated structure of 20Ne

### v1

@@ Abstract
We present the results obtained from the study of 20Ne--20Ne collisions at $\sqrt{s_{NN}}$= 6.8 TeV using the Monte Carlo Pythia8/Angantyr model. We construct a bi-pyramidal structure of 20Ne nucleus composed of five $\alpha$-clusters and compare its collision dynamics against those generated using Woods-Saxon nuclear density distribution. The results are further compared for different orientations of the 20Ne nuclei, including tip-tip, body-body, body-tip, and random orientations of the bi-pyramidal structure. The results show that the geometric arrangement of $\alpha$-clusters in the 20Ne nucleus significantly influences the particle production patterns in final state multiplicity, mean transverse momentum ($\langle p_{T}\rangle$), and elliptic flow coefficient ($v_{2}$). These results highlight the sensitivity of final state observables to the nuclear structure and orientation of colliding nuclei, providing insights into the dynamics of small collision systems, even in non-hydrodynamic framework.
@@ I Introduction
In recent years, the observation of collective behavior in small collision systems such as proton-proton (pp) and proton-nucleus (pA) collisions has challenged the traditional understanding of the conditions necessary for the emergence of quark-gluon plasma (QGP) and hydrodynamic-like phenomena [1, 2, 3, 4, 5, 6, 7]. The applicability of hydrodynamics to describe the collective behavior of the produced particles in small systems is quite challenging and has become a intriguing topic of investigation in heavy-ion physics. The initial geometry of the collision zone plays a crucial role in the hydrodynamic description of the strongly interacting matter produced in high energy nuclear collisions. There are several evidences in pp and pA collisions indicating fluctuations in the shape of proton from event to event  [8, 9, 10, 11] instead of spherical protons. Including a fluctuating nucleon substructure in model calculations improves the mean transverse momentum and harmonic flow coefficients description for charged and identified particles  [12]. The non-spherical nuclear shape either due to quadrapole deformation as seen in 238U and 129X or subnucleonic structure such as $\alpha$-clusters in 12C and 16O introduces anisotropies in the initial energy density profile which lead to pressure gradients and results in anisotropic expansion of the produced matter.
The effect of $\alpha$-cluster structure in C--C and O--O collisions on initial and final state observables has been studied extensively through various theoretical models [13, 14, 15] with harmonic flow coefficients as a measure sensitive to clusterization. The upcoming O--O data from the Large Hadron Collider (LHC) will validate and provide constraints on the theoretical predictions of the $\alpha$-cluster structure of the colliding nuclei. Like 16O, the 20Ne nucleus is also a prime candidate to study the influence of $\alpha$ clustering on the initial state geometry in nuclear collisions. The 20Ne nucleus is known to have a bi-pyramidal, balling-pin like structure composed of five $\alpha$-clusters with four $\alpha$-clusters arranged in a tetrahedral geometry which forms a configuration resembling a 16O nucleus, and the fifth $\alpha$-cluster is positioned along the symmetry axis,resulting in a 16O+$\alpha$ composite system. Several theoretical models have predicted the clustering in 20Ne, including the antisymmetrized quasicluster model (AQCM) [16], the antisymmetrized molecular dynamics (AMD)  [17], the hybrid-Brink-THSR wave function [18], the density functional theory (DFT) [19]. Some experimental searches have also been made which supports $\alpha$ clustering in 20Ne  [20, 21].
@@
These type of configurations has previously been investigated in collisions involving prolate shaped nuclei such as 238U and 129Xe, highlighting their impact on different final state observables [22, 23, 24, 25]. In these heavy nuclei, the deformation arises from the quadrupole deformation of the nuclear shape which leads to continuous and smooth spatial anisotropies. In contrast, $\alpha$ clustering in 20Ne introduces localized density hot-spots with discrete geometry arrangement, resulting in a more granular and fluctuating initial conditions. The presence of localized density fluctuations arising from alpha clustering can modify the initial eccentricities associated with different collision configurations of 20Ne, potentially leading to geometric features that differ from those observed in collisions involving smoothly deformed heavy ions.
In this Letter, we investigate the effect of the intrinsic structure of the 20Ne nucleus on the dynamics of small collision systems. We implement the $\alpha$-cluster structure of 20Ne in bi-pyramidal geometry in Monte Carlo Pythia8/Angantyr model. We primarily focus on some specific configurations of 20Ne--20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV. We aim to quantify how the elongated structure of 20Ne nuclei and its different orientations influences the particle production patterns in small collision systems.
II Framework
@@
Figure 2: Pseudorapidity distribution of charged hadrons in 20Ne--20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV for Woods-Saxon and bi-pyramidal geometries of 20Ne considering tip-tip, body-body, body-tip, and random orientations in (a) 0-5%, (b) 20-40%, and (c) 60-80% centrality classes. (d) $\langle dN_{ch}/d\eta\rangle$ as a function of centrality for different geometries and orientations.
To incorporate the $\alpha$-cluster structure of 20Ne in a bi-pyramidal geometry, we first constructed an oxygen nucleus based on $\alpha$ clustering, following the approach outlined in ref. [31]. In this configuration, all four $\alpha$ clusters are equidistant: three lie in the x-y plane, and the fourth is positioned along the negative z-axis. To form the 20Ne nucleus, a fifth $\alpha$ cluster is added along the positive z-axis at a slightly greater distance than the inter-cluster spacing used for the oxygen configuration. The distance is chosen to ensure that the resulting root-mean-square (rms) radius of the 20Ne nucleus closely matches the experimental value of 3.0055 fm [32]. In the bi-pyramidal configuration of 20Ne, the simulation yields an rms radius of 2.91 $\pm$ 0.09 fm. The position of the nucleons within the $\alpha$ clusters is sampled from the Woods-Saxon distribution represented using a three-parameter (3pF) Fermi distribution which is given by:
@@
where $\rho_{0}$ is the normalization constant, $R$ is the radius of the nucleus, $a$ is the surface thickness parameter, and $w$ is the skin thickness parameter. Each $\alpha$-cluster is assigned parameters $R$ = 0.964 fm, $a$ = 0.322 fm, and $w$ = 0.517 to define its nuclear charge density distribution [33]. Owing to the elongated structure of 20Ne, we have generated different orientations of the nucleus with respect to the beam direction. The specific orientations considered in the present article are tip-tip, body-body, body-tip, and random orientations. For generating random orientations, we randomize the positions of nucleons by performing three successive rotations on event by event basis for neon nucleus. For tip or body orientations, the long axis of the nucleus is aligned either along the beam direction or perpendicular to it, followed by a rotation in the x-y or y-z plane, respectively.
To study the effect of clustering, we have also simulated 20Ne collisions within the Angantyr model by employing the Woods-Saxon nuclear charge density distribution for the 20Ne nucleus as described by Eq. 1. The Woods-Saxon distribution parameters used are $R$ = 2.791 fm, $a$ = 0.698 fm, and $w$ = -0.168 [34], resulting in an rms radius of 2.99 $\pm$ 0.23 fm, which agrees well with the experimental value for the 20Ne nucleus. Fig. 1 shows the nucelon distribution of the 20Ne nucleus in bi-pyramidal geometry and Woods-Saxon distribution in x-y plane for 1 million events. Further, the distribution of nucleons y-z plane is also shown in the Fig. 1(c) after applying Euler rotations to the oxygen nucleus only. The resulting configuration, intended solely for visualization, reveals the balling-pin like structure of the 20Ne nucleus.
@@ III Results and Discussions
The results obtained for the pseudorapidity distribution of charged hadrons in 20Ne--20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV using the Pythia8/Angantyr model are displayed in Fig. 2. The pseudorapidity distribution is a crucial observable that provides insights into the particle production dynamics and the initial state geometry of the collision. The distribution is shown for the 0-5%, 20-40%, and 60-80% centrality classes, comparing the Woods-Saxon and bi-pyramidal geometries of 20Ne considering tip-tip, body-body, body-tip, and random orientations. $\langle dN_{ch}/d\eta\rangle$ as a function of centrality is shown in Fig. 2(d) for different geometries and orientations.
@@
Figure 3: (a) Transverse momentum ($p_{T}$) distribution of charged hadrons in 20Ne--20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV for Woods-Saxon and bi-pyramidal geometries of 20Ne considering tip-tip, body-body, body-tip, and random orientations in 0-5%, 20-40%, and 60-80% centrality classes. (b) Mean transverse momentum ($\langle p_{T}\rangle$) as a function of centrality for different geometries and orientations.
Figure 4: Directed flow ($v_{1}$), elliptic flow ($v_{2}$) and triangular flow ($v_{3}$) coefficients of charged hadrons as a function of transverse momentum ($p_{T}$) for Woods-Saxon and bi-pyramidal geometries of 20Ne considering tip-tip, body-body, body-tip, and random orientations in (a) 0-5%, (b) 20-40%, and (c) 60-80% centrality classes.
Once the orientation of the colliding nuclei is fixed, the situation changes drastically due to the change in the overlap area, which directly influences the number of participating nucleons. Among the various configurations, in central collisions, body-body collisions yield the highest multiplicity, followed by tip-tip, while body-tip results in the lowest. Both body-body and tip-tip configurations produce higher multiplicities than the Woods-Saxon baseline, and even body-tip is higher than random orientations. In body-body collisions, the transverse overlap between the spatially separated $\alpha$-clusters is maximized, resulting in a larger number of participating nucleons and consequently higher particle production. In tip-tip collisions, the overlap region is more compact with effective cluster interactions due to their alignment along the beam axis, leading to a slightly lower multiplicity than body-body collisions but still elevated multiplicity compared to Woods-Saxon geometry. In body-tip collisions, the disparity between the wide transverse profile of the body-oriented nucleus and the narrow shape of the tip-oriented nucleus results in a less effective overlap of $\alpha$-clusters, thereby reducing the number of nucleon-nucleon interactions. As a result, the body-tip collisions yield the lowest multiplicity among the fixed orientations, although it is still higher than the random orientation case, where cluster alignment is not optimized. As collisions become non-central, the overlap region decreases, leading to misalignment of $\alpha$-clusters across different orientations. This weakens the geometrical advantages seen in central collisions, causing significant changes in multiplicity among fixed orientations. The multiplicity density of body-tip collisions closely follows the trend of random orientations observed in different centrality classes, while tip-tip collisions show higher multiplicity in central and mid-central collisions than the random orientation, but as the collisions become more peripheral, the multiplicity density becomes lower than that of random orientations.
@@
Fig. 4 shows the directed flow ($v_{1}$), elliptic flow ($v_{2}$) and triangular flow ($v_{3}$) coefficients of charged hadrons as a function of transverse momentum ($p_{T}$) for Woods-Saxon and bi-pyramidal geometries of 20Ne considering tip-tip, body-body, body-tip, and random orientations in three centrality intervals. The directed and triangular flow coefficients are negligible in all the configurations and centrality intervals, while the elliptic flow coefficients show a significant dependence on the collision geometry and centrality. In principle, the momentum conservation, nuclear configurations, fluctuations in nucleon positions and nuclear clustering should influence the directed and triangular flow coefficients. Although, the presence of $\alpha$-clustering in 20Ne--20Ne collisions introduces geometric imbalances, the pressure gradients necessary to convert these into net longitudinal momentum flow is absent in Angantyr model. As a result, the $\alpha$-clustering and different geometrical orientations of 20Ne nuclei may effect multiplicity or pseudorapidity distribution of charged hadrons, but they are not sufficient to generate significant directed flow. Similarly, the triangular flow coefficient arises from the fluctuations in the initial geometry of the collisions which are already implemented in the Angantyr model and especially in $\alpha$-clustered geometry of 20Ne. Despite that, the collective respose mechanism necessary to translate these fluctuations into a measurable triangular flow is not present in Angantyr model, leading to negligible triangular flow coefficients.
The elliptic flow coefficients, however, show a significant dependence on the collision geometry and centrality. The elliptic flow generally arises from the initial geometrical anisotropy of the nuclear overlap region, which is influenced by the shape and orientation of the colliding nuclei. 20Ne--20Ne collisions with $\alpha$-clusters enhances local density variations and anisotropies depending on the orientation of the collision leading to an azimuthally asymmetric distribution of NN collisions which results in a directionally biased particle production. The localized density of NN collisions regulates MPI and string formation, regions with greater overlap experience more MPI producing higher string densities. Hadronization of these strings are modeled through Lund string fragmentation model, which is sensitive to the local density of strings. Anisotropy in the string configuration can be reflected in the angular distribution of the final state hadrons, leading to the emergence of elliptic flow even in the absence of collectivity. Additionally, resonance decays can also modify these angular correlations, further contributing to the observed elliptic flow signal.
@@ IV Conclusions
In conclusion, we have investigated the effect of the intrinsic structure of the 20Ne nucleus on the dynamics of small collision systems using the Pythia8/Angantyr model. By implementing the $\alpha$-cluster structure of 20Ne in a bi-pyramidal geometry, we have explored the influence of different orientations of the elongated structure on the particle production patterns in 20Ne--20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV. Our results show that the pseudorapidity distribution of charged hadrons is significantly modified by the $\alpha$-cluster structure, with the Woods-Saxon distribution yielding higher particle yields in central collisions compared to the bi-pyramidal geometry in random orientation. The different orientations of the 20Ne nuclei lead to distinct particle production patterns, with body-body collisions resulting in the highest multiplicity in central collisions, while body-tip collisions yield the lowest. The transverse momentum distribution of charged hadrons shows only slight variations between Woods-Saxon and bi-pyramidal geometries, indicating that the $\alpha$-cluster structure does not significantly affect the $p_{T}$ distribution. The mean transverse momentum of charged hadrons decreases with increasing centrality for all configurations, except for body-body collisions, which exhibit an increasing trend in mid-central collisions. The elliptic flow coefficients show a significant dependence on the collision geometry and centrality, reflecting the influence of the initial geometrical anisotropy of the nuclear overlap region. However, the directed and triangular flow coefficients remain negligible across all configurations and centrality intervals, indicating that the presence of $\alpha$-clustering and different orientations of 20Ne nuclei are not sufficient to generate significant directed and triangular flow in the Angantyr model.

### v2

@@ Abstract
We present the results obtained from the study of 20Ne-20Ne collisions at $\sqrt{s_{NN}}$= 6.8 TeV using the Monte Carlo Pythia8/Angantyr model. Angantyr models heavy-ion collisions as a superposition of independent nucleon-nucleon (NN) collisions, without incorporating collective effects. We construct a bi-pyramidal structure of 20Ne nucleus composed of five $\alpha$-clusters and compare its collision dynamics against those generated using Woods-Saxon nuclear density distribution. The results are further compared for different orientations of the 20Ne nuclei, including tip-tip, body-body, body-tip, and random orientations of the bi-pyramidal structure. The results show that the geometric arrangement of $\alpha$-clusters in the 20Ne nucleus significantly influences the particle production patterns in final state particle multiplicity and mean transverse momentum ($\langle p_{T}\rangle$). These results highlight the sensitivity of final state observables to the nuclear structure and orientation of colliding nuclei, providing insights into the dynamics of small collision systems, even in non-hydrodynamic framework.
@@ I Introduction
In recent years, the observation of collective behavior in small collision systems such as proton-proton (pp) and proton-nucleus (pA) collisions has challenged the traditional understanding of the conditions necessary for the emergence of quark-gluon plasma (QGP) and hydrodynamic-like phenomena [1, 2, 3, 4, 5, 6, 7]. The applicability of hydrodynamics to describe the collective behavior of the produced particles in small systems is quite challenging and has become an intriguing topic of investigation in heavy-ion physics. The initial geometry of the collision zone plays a crucial role in the hydrodynamic description of the strongly interacting matter produced in high energy nuclear collisions. There are several evidences in pp and pA collisions indicating fluctuations in the shape of proton from event to event  [8, 9, 10, 11]. The inclusion of a fluctuating nucleon substructure in model calculations seemed to improve the mean transverse momentum and harmonic flow coefficients description for charged and identified particles  [12]. The non-spherical nuclear shape either due to quadrupole deformation as seen in 238U and 129Xe or subnucleonic structure such as $\alpha$-clusters in 12C and 16O introduce anisotropies in the initial energy density profile which possibly leads to collective expansion of the produced matter due to formation of pressure gradients.
[…]
@@
These type of configurations has been investigated previously in collisions involving prolate shaped nuclei such as 238U and 129Xe, highlighting their impact on different final state observables [22, 23, 24, 25]. In these heavy nuclei, the deformation arises from the quadrupole deformation of the nuclear shape which leads to continuous and smooth spatial anisotropies. In contrast, $\alpha$ clustering in 20Ne introduces localized density hot-spots with discrete geometry arrangement, resulting in a more granular and fluctuating initial conditions. The presence of localized density fluctuations arising from alpha clustering can modify the initial eccentricities associated with different collision configurations of 20Ne, potentially leading to geometric features that differ from those observed in collisions involving smoothly deformed heavy ions.
In this Letter, we investigate the effect of the intrinsic structure of the 20Ne nucleus on the dynamics of small collision systems. We implement the $\alpha$-cluster structure of 20Ne in bi-pyramidal geometry in Monte Carlo Pythia8/Angantyr model. We primarily focus on some specific configurations of 20Ne-20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV. We aim to quantify how the elongated structure of 20Ne nuclei and its different orientations influences the particle production patterns in small collision systems, focusing specifically on the absence of hydrodynamic evolution. This choice allows us to isolate the direct impact of the intrinsic nuclear geometry on particle production in small systems without medium-induced effects.
II Collision Framework
@@
Figure 2: Pseudorapidity distribution of charged hadrons in 20Ne-20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV for Woods-Saxon and bi-pyramidal geometries of 20Ne considering tip-tip, body-body, body-tip, and random orientations in (a) 0-5%, (b) 20-40%, and (c) 60-80% centrality classes. (d) $\langle dN_{ch}/d\eta\rangle$ as a function of centrality for different geometries and orientations.
To incorporate the $\alpha$-cluster structure of 20Ne in a bi-pyramidal geometry, an oxygen nucleus based on $\alpha$ clustering was constructed following the approach outlined in ref. [31]. In this configuration, all four $\alpha$ clusters are equidistant: three lie in the x-y plane, and the fourth is positioned along the negative z-axis. To form the 20Ne nucleus, a fifth $\alpha$ cluster is added along the positive z-axis at a slightly greater distance than the inter-cluster spacing used for the oxygen configuration. The distance is chosen to ensure that the resulting root-mean-square (rms) radius of the 20Ne nucleus closely aligns with the experimental value of 3.0055 fm [32]. In the bi-pyramidal configuration of 20Ne, the simulation yields an rms radius of 2.91 $\pm$ 0.09 fm. The position of the nucleons within the $\alpha$ clusters is sampled from the Woods-Saxon distribution represented using a three-parameter (3pF) Fermi distribution which is given by:
@@
where $\rho_{0}$ is the normalization constant, $R$ is the radius of the nucleus, $a$ is the surface thickness parameter, and $w$ is the skin thickness parameter. Each $\alpha$-cluster is assigned parameters $R$ = 0.964 fm, $a$ = 0.322 fm, and $w$ = 0.517 to define its nuclear charge density distribution [33]. Owing to the elongated structure of 20Ne, different orientations of the nucleus with respect to the beam direction were generated. The specific orientations considered in the present article are tip-tip, body-body, body-tip, and random orientations. For generating random orientations, we randomize the positions of nucleons by performing three successive rotations on an event by event basis for neon nucleus. For tip or body orientations, the long axis of the nucleus is aligned either along the beam direction or perpendicular to it, followed by a rotation in the x-y or y-z plane, respectively.
To study the effect of clustering, we have also simulated 20Ne collisions within the Angantyr model by employing the Woods-Saxon nuclear charge density distribution for the 20Ne nucleus as described by Eq. 1. The Woods-Saxon distribution parameters used are $R$ = 2.791 fm, $a$ = 0.698 fm, and $w$ = -0.168 [34], resulting in an rms radius of 2.99 $\pm$ 0.23 fm, which agrees well with the experimental value for the 20Ne nucleus. Fig. 1 shows the nucelon distribution of the 20Ne nucleus in bi-pyramidal geometry and Woods-Saxon distribution in x-y plane for one million events. Further, the distribution of nucleons y-z plane is also shown in the Fig. 1(c) after applying Euler rotations to the oxygen nucleus only. The resulting configuration, intended solely for visualization, reveals the bowling-pin like structure of the 20Ne nucleus.
@@ III Results and Discussions
The results obtained for the pseudorapidity distribution of charged hadrons in 20Ne-20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV using the Pythia8/Angantyr model are displayed in Fig. 2. The pseudorapidity distribution is a crucial observable that provides insights into the particle production dynamics and the initial state geometry of the collision. The distribution is shown for the 0-5%, 20-40%, and 60-80% centrality classes, comparing the Woods-Saxon and bi-pyramidal geometries of 20Ne considering tip-tip, body-body, body-tip, and random orientations. $\langle dN_{ch}/d\eta\rangle$ as a function of centrality is shown in Fig. 2(d) for different geometries and orientations.
@@
Figure 3: (a) Transverse momentum ($p_{T}$) distribution of charged hadrons in 20Ne-20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV for Woods-Saxon and bi-pyramidal geometries of 20Ne considering tip-tip, body-body, body-tip, and random orientations in 0-5%, 20-40%, and 60-80% centrality classes. (b) Mean transverse momentum ($\langle p_{T}\rangle$) as a function of centrality for different geometries and orientations.
Once the orientation of the colliding nuclei is fixed, the situation changes drastically due to the change in the overlap area, which directly influence the number of participating nucleons. Among the various configurations, in central collisions, body-body collisions yield the highest multiplicity, followed by tip-tip, while body-tip results in the lowest. Both body-body and tip-tip configurations produce higher multiplicities than the Woods-Saxon baseline, and even body-tip is higher than random orientations. In body-body collisions, the transverse overlap between the spatially separated $\alpha$-clusters is maximized, resulting in a larger number of participating nucleons and consequently higher particle production. In tip-tip collisions, the overlap region is more compact with effective cluster interactions due to their alignment along the beam axis, leading to a slightly lower multiplicity than body-body collisions but still elevated multiplicity compared to Woods-Saxon geometry. In body-tip collisions, the disparity between the wide transverse profile of the body-oriented nucleus and the narrow shape of the tip-oriented nucleus results in a less effective overlap of $\alpha$-clusters, thereby reducing the number of nucleon-nucleon interactions. As a result, the body-tip collisions yield the lowest multiplicity among the fixed orientations, although it is still higher than the random orientation case, where cluster alignment is not optimized. As collisions become non-central, the overlap region decreases, leading to misalignment of $\alpha$-clusters across different orientations. This weakens the geometrical advantages seen in central collisions, causing significant changes in multiplicity among fixed orientations. The multiplicity density of body-tip collisions closely follows the trend of random orientations observed in different centrality classes, while tip-tip collisions show higher multiplicity in central and mid-central collisions than the random orientation, but as the collisions become more peripheral, the multiplicity density becomes lower than that of random orientations.
@@ IV Conclusions
In conclusion, we have investigated the effect of the intrinsic structure of the 20Ne nucleus on the dynamics of small collision systems using the Pythia8/Angantyr model. By implementing the $\alpha$-cluster structure of 20Ne in a bi-pyramidal geometry, we have explored the influence of different orientations of the elongated structure on the particle production patterns in 20Ne-20Ne collisions at $\sqrt{s_{NN}}$ = 6.8 TeV. Our results show that the pseudorapidity distribution of charged hadrons is significantly modified by the $\alpha$-cluster structure, with the Woods-Saxon distribution yielding higher particle yields in central collisions compared to the bi-pyramidal geometry in random orientation. The different orientations of the 20Ne nuclei lead to distinct particle production patterns, with body-body collisions resulting in the highest multiplicity in central collisions, while body-tip collisions yield the lowest. The transverse momentum distribution of charged hadrons shows only slight variations between Woods-Saxon and bi-pyramidal geometries, indicating that the $\alpha$-cluster structure does not significantly affect the $p_{T}$ distribution. The mean transverse momentum of charged hadrons decreases with increasing centrality for all configurations, except for body-body collisions, which exhibit an increasing trend in mid-central collisions.
The findings highlight the sensitivity of small-system dynamics to the underlying nuclear geometry, demonstrating that the intrinsic $\alpha$-cluster structure of light nuclei such as 20Ne leaves observable imprints on particle production, even in the absence of hydrodynamic evolution. It will also be interesting to investigate the role of different orientations in final-state observables within a hydrodynamic framework, so that forthcoming experimental results can be compared to scenarios with and without medium effects, offering valuable insight into the interplay between nuclear structure and collective dynamics.